Can someone explain the equations I put in the box in detail?

12 dB dt 4π r. 4T r 2π r

2 0
Answer #1

\underset{F}{\rightarrow} = q (\underset{v}{\rightarrow}x \underset{B}{\rightarrow} )

The above equation is used to determine the magnetic force experienced by a charge particle having charge "q" on it and moving with velocity "v" in a region having magnetic field "B"

\underset{F}{\rightarrow} = i \underset{l}{\rightarrow} x \underset{B}{\rightarrow}

The above equation is used to determine the magnetic force experienced by a current carrying wire of length "l" and carrying current "i" in a region having magnetic field "B".

\underset{B}{\rightarrow}= (\mu _{o}/4\pi) (q \underset{v}{\rightarrow} x \hat{r})/r2

The above equation gives the magnetic field produced by a moving charge having charge "q" and velocity "v" , at a distance "r" from the location where the magnetic field is to be determined.

\underset{B}{\rightarrow}= (\mu _{o}/2\pi) (I/r)

The above equation gives the magnetic field produced by a long wire carrying current "I" at a distance "r" from the wire.

\phi _{m} = \int \underset{B}{\rightarrow}.\underset{dA}{\rightarrow}

The above equation gives the magnetic flux produced by magnetic field through a surface area.

E = - d\phi/dt

This equation gives the induced emf